### v2

@@ Appendix A Graphs of the $B$-index
Figure A.1: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the LR classification.
@@
Figure A.2: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the SVM classification.
@@
Figure A.3: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the GB classification.
@@
Figure A.4: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the MLP classification.
@@
Figure A.5: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the RF classification.

### v3

@@ Appendix B The $B$-index and samples sizes
Figure B.1: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the LR classification.
@@
Figure B.2: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the SVM classification.
@@
Figure B.3: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the GB classification.
@@
Figure B.4: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the MLP classification.
@@
Figure B.5: The plots of $B_{n,p}$ (vertical axis) for various sample sizes $n$ (horizontal axis) and parameter $p$ values when using the RF classification.